OpenStudy (anonymous):

Could someone help me with: Suppose sin(β) = 1/7 Find the exact value of sec(β) for 0 < β < pi/2 Thanks in advance!

OpenStudy (ash2326):

@Rhope we know sec x = 1/cos x could you find cos x?

OpenStudy (cwrw238):

my draw function isnt working so i'll need to do this another way relation between sin B and cos B is cos^2 B = 1 - sin^2 B = 1 - (1/7)^2 so cos B = sqrt(1 - 1/49) = sqrt 48/49 now sec B = 1/ cos B so sec B =sqrt ( 49 / 48) which we can simplify further

OpenStudy (cwrw238):

sec B = 7 / sqrt 48 = 7 / 4sqrt3 = 7sqrt3 / 4

OpenStudy (anonymous):

I figured it out using the triangle and SOHCAHTOA. Thanks though everyone! (:

OpenStudy (cwrw238):

good - thats the way i was going to show you if my draw thing was working!